OpenStudy (anonymous):

y - 1 = -2 ( x + 3 ) 2x + 3y = 6 How do I solve and graph this out? Please help? :(((

OpenStudy (anonymous):

y = -2(x +3) +1 y = -2x -3 +1 y = -2x -2 Plug in this value of y in 2x +3y = 6 2x +3*(-2x -2) = 6 2x -6x -6 = 6 -4x = 12 x = -3 Plug in x and find y. x =-3, y = -2x -2 = -2*(-3) -2 = 6 -2 = 4 Answer (-3,4)

OpenStudy (anonymous):

the first one is in point slope and the second one is in standard form first one is easy to graph. you have a point built in, and the slope, the point would be (-3, 1) because the y value is 1 and x is -3 according to the equation slope of -2 as well the second equation you will have to change to slope intercept 2x+3y=6 3y=-2x+6 y= (-2/3)x+2 plot points and graph